OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (–7, 9); m = 4/5

OpenStudy (anonymous):

y + 7 = 4/5(x – 9) y – 7 = 4/5(x + 9) y – 9 = 4/5(x + 7) y + 9 = 4/5(x – 7)

OpenStudy (anonymous):

...And lol if you have time could you walk me through this question as well? The table shows the height of a tree as it grows. What equation in slope-intercept form gives the tree's height at any time? Time Height (months) (inches) 2 14 4 23 6 32 8 41 y = 9x + 2 y = 9/2x + 5 y = 9/2x y = 5x + 9/2

OpenStudy (anonymous):

Do you know point slope form?

OpenStudy (anonymous):

Not really :(

OpenStudy (anonymous):

Sorry. I hate math and I suck at it lol

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\] For m, plug in the slope. For y1, plug in the y coordinate (9), and for x1, plug in the x coordinate (-7)

OpenStudy (anonymous):

Let me know what you get when you plug it all in

OpenStudy (anonymous):

I'm still kind of confused.. I'm so sorry

OpenStudy (anonymous):

Do you see the equation i posted?

OpenStudy (anonymous):

y - 9 = m(x-7)

OpenStudy (anonymous):

the first part is perfect, but it's not supposed to be -7. Its a double negative, which makes it a +7